# Improving threshold for fault-tolerant color code quantum computing by flagged weight optimization

### v1

@@ V Conclusion
In this work, we have proposed the flagged weight optimization, a decoder weight optimization method using flag qubits. Utilizing flag values allows us to set more optimal weights, leading to more accurate decoding. Also, the flag gadget reduces the depth of the syndrome measurement circuit and thus suppresses the impact of errors. By applying this method to the integer programming decoder, we improved the circuit-level threshold of the (4.8.8) color code from the existing 0.14% to around 0.27%. In the case of the (6.6.6) color code, we achieved the circuit-level threshold of around 0.36%, which is nearly identical to the highest value obtained in the previous works that employ the same noise model. In both cases, the logical error rates at the low physical error rates were improved by almost one order of magnitude compared to the single ancilla method. By utilizing this approach, it is expected that color code-based FTQC, which enables the transversal implementation of all logical Clifford gates, will become more promising. This method can be applied to other weight-based decoders. One can also use this method to improve the threshold of wider classes of QECCs, such as high-rate quantum LDPC codes, which have high-weight stabilizer generators.
@@ Acknowledgements.
This work is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant No. JPMXS0118067394 and JPMXS0120319794, JST COI-NEXT Grant No. JPMJPF2014, and JST Moonshot R&D Grant No. JPMJMS2061.
@@ Appendix A Details of weight
We describe the derivation of the decoder weight defined by Eqs. (12) and (13). For simplicity, we assume that only data errors occur independently, with each error not having identical probabilities. In this assumption, the probability of a certain error event $E$ occurring is given by
@@
where $e_{i}$ denotes the $i$-th data error. The task of decoding is to estimate the most likely error given the syndrome $S$; that is to estimate
@@
Thus, under the assumption that error probabilities are not identical across all data qubits, the task of decoding is to estimate the errors satisfying the syndrome constraints and minimizing the sum of the following values:
@@
Therefore, by setting the weights of the decoder to the values defined by Eq. (25), it is possible to decode taking into account the difference in each error probability. Weights for measurement errors can also be derived through the same reasoning if we assume that the measurement errors occur independently.
Figure 18: A typical example of performance improvement by the deflagging procedure. (a) Without deflagging. When errors occur in the syndrome measurement circuit and cause the measurement outcomes shown in the left figure, decoding based on the conditional error probabilities in Fig. 17(a) leads to the estimation of errors as depicted in the right figure. It results in a logical error after the recovery operation. The actual data errors are also displayed in the left figure, even though they cannot be identified in actual scenarios. (b) With deflagging. After performing the deflagging procedure and then decoding based on the conditional error probabilities in Fig. 17(b), errors are estimated as illustrated in the rightmost figure, leading to a successful error correction.
@@ Appendix B Details of deflagging
Here, details of the deflagging procedure performed in this work are described. In the two-qubit flag gadget, we employ a slightly different way of applying Pauli operators from Refs. [32, 33]. In Refs. [32, 33], when a flag is triggered, a Pauli operator is applied to one of the data qubits connected to the flag qubit. In this study, when a flag is triggered in the two-qubit flag gadget, a proper type of Pauli operator is applied to all data qubits connected to the syndrome qubit. When performing flagged weight optimization in addition to the deflagging procedure, we verified in the numerical simulation that the performance of both is almost the same, so either method can be used. In the four-qubit flag gadget, if all three flags are triggered, a proper type of Pauli operator is applied to all data qubits connected to the top flag qubit and the syndrome qubit in Fig. 8(b).
@@
We show a typical error event that can be corrected by performing a deflagging procedure in addition to the flagged weight optimization. We provide an example using the distance-5 (4.8.8) color code when the physical error rate is $p=10^{-4}$. Let us consider a situation where errors occur in only one time step and not in other time steps to show a typical example. Here, the weights of measurement errors are not shown for simplicity. The conditional data error probabilities estimated in our numerical simulation when certain two flags are triggered are shown in Fig. 17. Figs. 17(a) and 17(b) present the conditional data error probabilities without and with the deflagging procedure, respectively. Note that these values are not always estimated due to statistical errors. Suppose some errors occurred in the syndrome measurement circuit, resulting in the measurement outcomes shown on the left in Fig. 18(a). The actual data errors are also shown, although they are unknown in real situations. When decoding is performed with the weights determined by the conditional error probabilities of Fig. 17(a), an error event on the right of Fig. 18(a) is estimated. Performing a recovery operation based on this decoding result leads to a logical error. On the other hand, when the deflagging procedure is performed for this error event, it becomes the central figure in Fig. 18(b). Decoding with the weights determined by the conditional error probabilities in Fig. 17(b) estimates an error event on the rightmost figure of Fig. 18(b). This leads to a successful error correction. What we have explained above is just one typical example, so we compare the logical error rates to verify if the deflagging procedure improves the performance. Fig. 19 shows the logical $Z$ error rates for the (4.8.8) color code in cases where the deflagging procedure is performed and not performed. Fig. 19 indicates that the deflagging procedure improves the logical error rates.

### v2

@@ V Conclusion
In this work, we have proposed flagged weight optimization (FWO), a decoder weight optimization method using conditional error probabilities conditioned on the measurement outcomes of flag qubits. Utilizing flag values allows us to set more optimized weights, leading to more accurate decoding. Also, the cat states reduce the depth of the syndrome measurement circuit and thus suppresses the impact of errors, as also noted in Ref. [24]. By applying this method to the integer programming decoder, we improved the circuit-level threshold of the (4.8.8) color code from the existing 0.14% to around 0.27%. In the case of the (6.6.6) color code, we achieved the circuit-level threshold of around 0.36%, which is identical within statistical errors to the highest value of 0.37(1)obtained in the previous works that employ the same noise model. In both cases, an effective code distance is also improved compared to the single ancilla method, meaning that FWO helps correct large hook errors that arise from relatively few faults. Thereby, the achieved logical error rates at low physical error rates are almost one order of magnitude lower than the single ancilla method with the same code distance. We note that the threshold values obtained here are calculated by using the code distances up to $d=9$. A numerical experiment at larger code distances is very important, but it is left for future work.
We also verified, even when comparing to the single ancilla method with a code distance higher than our method but requires the similar number of qubits, our method achieves lower logical error rates in most cases. By utilizing this approach, it is expected that color code-based FTQC, which enables the transversal implementation of all logical Clifford gates, will become more promising. This method can be applied to other weight-based decoders. One can also use this method to improve the threshold of wider classes of QECCs, such as high-rate quantum LDPC codes, which have high-weight stabilizer generators.
Regarding the use of cat states, cat states offer benefits such as reducing the connectivity requirements for hardware, increasing the effective code distance, and decreasing the circuit depth. However, they also increase the number of qubits and circuit width. Considering the increased number of qubits, it is generally an open question as to whether using cat states for measuring syndrome is worth it, purely from the perspective of improving logical error rates. Nonetheless, we verified that they are worth it for improving logical error rates when performing FWO and deflagging, even taking into account the increased number of qubits, as discussed in Sec. IV.
In this work, we used flag information to optimize the edge weights in the decoding graph. On the other hand, one can use flag information as additional parity checks by adding nodes corresponding to each flag qubit and edges connecting these nodes to other relevant nodes in the decoding graph. That method could achieve lower logical error rates than FWO, but FWO has an advantage over it. If we add nodes corresponding to each flag qubit and relevant edges to the decoding graph, the computational complexity of the decoding increases considerably, and more importantly, the decoding algorithm no longer works in certain decoders due to being unable to handle the additional nodes and edges. On the other hand, in FWO, we do not add additional nodes and edges to the decoding graph. Hence, if a decoder works in the phenomenological noise model, we can use FWO with that decoder in the circuit-level noise model. We believe that this advantage is important, because a decoder that is more promising than existing decoders in terms of decoding time or accuracy, but with limitations on the decoding graph, may be developed in the future. Therefore, a versatile method that is applicable to a wider variety of decoders is useful.
@@ Acknowledgements.
The authors thank Theerapat Tansuwannont for helpful discussions. This work is supported by MEXT Quantum Leap Flagship Program (MEXT Q-LEAP) Grant No. JPMXS0118067394 and JPMXS0120319794, JST COI-NEXT Grant No. JPMJPF2014, and JST Moonshot R&D Grant No. JPMJMS2061.
@@ Appendix A Details of weight
We describe the derivation of the decoder weight defined by Eqs. (9) and (10) [30]. We consider the situation where data errors and measurement errors occur independently, with each error not having identical probabilities. In this situation, the probability of a certain error event $E$ occurring is given by
@@
where $e_{i}$ denotes the $i$-th data error or measurement error. The task of decoding is to estimate the most likely error given the syndrome $S$; that is to estimate
@@
Thus, under the assumption that error probabilities are not identical across all data errors and measurement errors, the task of decoding is to estimate the errors satisfying the syndrome constraints and minimizing the sum of the following values:
@@
Therefore, by setting the weights of the decoder to the values defined by Eq. (25), it is possible to decode taking into account the difference in each error probability.
Figure 18: A typical example of performance improvement by the deflagging procedure in the case of $X$ error correction. In this figure, data errors are $X$ type, and the syndromes and flags are the ones that indicate the occurrence of $X$ errors. (a) Without deflagging. When errors occur in the syndrome measurement circuit and cause the measurement outcomes shown in the left figure, decoding based on the conditional error probabilities in Fig. 17(a) leads to the estimation of errors as depicted in the right figure. It results in a logical error after the recovery operation. The actual data errors are also displayed in the left figure, even though they cannot be identified in actual scenarios. (b) With deflagging. After performing the deflagging procedure and then decoding based on the conditional error probabilities in Fig. 17(b), errors are estimated as illustrated in the rightmost figure, leading to a successful error correction.
@@ Appendix B Details of deflagging
Here, details of the deflagging procedure performed in this work are described. In the two-qubit flag gadget, we employ a slightly different way of applying Pauli operators from Refs. [38, 39]. In Refs. [38, 39], when a flag is triggered, a Pauli operator is applied to one of the data qubits connected to the flag qubit. In this study, when a flag is triggered in the two-qubit flag gadget, a proper type of Pauli operator is applied to all data qubits connected to the syndrome qubit. Here, proper type means $X$ type for the flags triggered during the $X$ stabilizer measurement and $Z$ type for those triggered during the $Z$ stabilizer measurement. When performing FWO in addition to the deflagging procedure, we verified that the improvement in logical error rates by our deflagging procedure is the same within the range of error bars as that when we perform the deflagging procedure proposed in Refs. [38, 39]. Thus, either way of deflagging procedure can be used. In the four-qubit flag gadget, if all three flags are triggered, a proper type of Pauli operator is applied to all data qubits connected to the top flag qubit and the syndrome qubit in Fig. 8(b).
@@
We show a typical error event that can be corrected by performing a deflagging procedure in addition to FWO. We provide an example using the distance-5 (4.8.8) color code when the physical error rate is $p=10^{-4}$. Let us consider a situation where errors occur in only one time step and not in other time steps to show a typical example. Here, the weights of edges corresponding to measurement errors are not shown for simplicity. We consider the case of $X$ error correction; that is, we focus on $X$ data errors, and the syndromes and flags considered below indicate the occurrence of $X$ errors. The conditional data error probabilities estimated in our numerical simulation when certain two flags are triggered are shown in Fig. 17. Figs. 17(a) and 17(b) present the conditional data error probabilities without and with the deflagging procedure, respectively. Note that estimated conditional probabilities have statistical fluctuations depending on the number of samples and the probabilities that the set of flags is triggered. Suppose some errors occurred in the syndrome measurement circuit, resulting in the measurement outcomes shown on the left in Fig. 18(a). The actual data errors are also shown, although they are unknown in real situations. When decoding is performed with the weights determined by the conditional error probabilities of Fig. 17(a), an error event on the right of Fig. 18(a) is estimated. Performing a recovery operation based on this decoding result leads to a logical error. On the other hand, when the deflagging procedure is performed for this error event, it becomes the central figure in Fig. 18(b). Note that the Pauli operators applied in the deflagging procedure here are $X$ type, because we are considering the case of $X$ error correction in this example. Decoding with the weights determined by the conditional error probabilities in Fig. 17(b) estimates an error event on the rightmost figure of Fig. 18(b). This leads to a successful error correction. What we have explained above is just one typical example, so we compare the logical error rates to verify if the deflagging procedure improves the performance. Fig. 19 shows the logical $Z$ error rates for the (4.8.8) color code in cases where the deflagging procedure is performed and not performed. Fig. 19 indicates that the deflagging procedure improves the logical error rates.
Appendix C Analysis of the improvement achieved by FWO
The numerical results in Sec. IV are the results of the combined contributions from several techniques described in Sec. III, namely the use of cat states, FWO, and deflagging. For that reason, it is unclear how much the FWO, the main proposed technique, contributes to the results. Thus, we here calculate the logical error rates when using cat states and performing deflagging but without FWO, and compare with the results obtained in Sec. IV to clarify how much FWO itself improves the performance. Logical $Z$ error rates for the (4.8.8) color code without performing FWO, but employing cat states and deflagging are shown in the dotted lines of Fig. 20. The solid lines in Fig. 20 and the solid lines in Fig. 15(b) are the same data. We also show the values obtained by fitting the logical error rates of the dotted lines in Fig. 20 with Eqs. (16) and (17) in Table 5. From Fig. 20, it can be seen that the logical error rates considerably decrease by performing FWO. Also, by comparing Tables 5 and 2, we can see that $p_{\mathrm{th}}^{*}$, $p_{\mathrm{th}}^{\times}$, and the effective code distance are all greatly improved by FWO. These results indicate that FWO itself contributes significantly to the improvements observed in Sec. IV.
Figure 20: Logical $Z$ error rates of the (4.8.8) color code with and without FWO. Obtained scaling threshold and cross threshold for the data of the dotted lines are $p_{\mathrm{th}}^{*}=0.183(2)\%$ and $p_{\mathrm{th}}^{\times}=0.1843(9)\%$, respectively.
Table 5: Fitting parameters for the logical $Z$ error rates of the (4.8.8) color code without FWO.
$c$ $\alpha$ $p_{\mathrm{th}}^{*}$ $p_{\mathrm{th}}^{\times}$ Without FWO 0.090(3) 0.583(3) 0.00183(2) 0.001843(9)